OpenStudy (anonymous):

What is F(-2)? If your answer is not a whole number, enter it as a fraction in lowest terms,using the slash mark as a fraction bar F(x)= x^3+4/ x+1

OpenStudy (anonymous):

@Callisto will you please help me?

OpenStudy (anonymous):

@Directrix could you help please?

OpenStudy (anonymous):

@Zarkon could you help me please?

Directrix (directrix):

F(x)= (x^3+4)/ (x+1) Is F(x) supposed to be the rational function above. You did not show any grouping symbols so I am not sure. F(x)= x^3+4/ x+1 is not the same function as F(x)= (x^3+4)/ (x+1). So we need to agree on what F(x) is as a function before we calculate anything.